# Zeroth-order log-concave sampling

### v1

@@ 2.1 Mixing analysis
The data-processing inequality (DPI) is a handy tool in information theory; for any probability measures $\mu,\nu$, Markov kernel $P$, $\varphi$-divergence $D_{\varphi}$, and $q\in(1,\infty]$, it holds that
@@
However, this is not sufficient to establish a mixing rate. It turns out that adding an independent Gaussian random variable to $X_{k}$ (i.e., convolving $\operatorname{law}X_{k}$ with $\gamma_{h}$) makes $\operatorname{law}X_{k}$ contract directly toward $\pi^{X}*\gamma_{h}=\pi^{Y}$, with the contraction rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$.
@@
where $\nu_{t}:=\nu*\gamma_{t}$ denotes the probability measure obtained by convolving $\nu$ with $\gamma_{t}$. Clearly, $0\leq\eta_{D}(\mu,t)\leq 1$ by the DPI.
Bounding $\eta_{D}$ relies on the well-known de Bruijn identity, $\partial_{t}\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{1}{2}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$, where the relative Fisher information is defined as $\mathsf{FI}(\nu\mathbin{\|}\mu):=\mathbb{E}_{\nu}[|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]$; see rigorous proofs by Barron [Bar84] and Klartag and Ordentlich [KO25] (for its generalizations). Other derivations can be found in [Guo09, VW23, CCSW22], assuming that probability measures involved in proofs are sufficiently smooth to justify differentiation under the integral sign and vanishment of boundary terms in integration by parts.
Combining this identity with, for example, ($\mathsf{LSI}$) for $\mu$ (i.e., $\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})\leq\nicefrac{{C_{\mathsf{LSI}}(\mu_{t})}}{{2}}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$) and $C_{\mathsf{LSI}}(\mu_{t})\leq C_{\mathsf{LSI}}(\mu)+t$ [Cha04], one can obtain a differential inequality in terms of $\mathsf{KL}$, and solving this leads to
@@
Under similar regularity assumptions, the SDPI for Rényi divergence under ($\mathsf{PI}$), ($\mathsf{LSI}$), and Latała–Oleszkiewicz inequality was studied in [CCSW22] and further extended to constrained probability measures in [KVZ24]. These results were rigorously unified and investigated in [KO25] under a $\varphi$-Sobolev inequality [Cha04] (i.e., a generalization of ($\mathsf{PI}$) and ($\mathsf{LSI}$)), with regularity assumptions explicitly identified. We refer readers to [KO25, Regularity conditions] for a detailed account.
Specifically, [CCSW22] showed that under the regularity assumptions, for any $q\geq 2$ and $\nu\ll\mu$, the following bounds hold:
$\mathcal{R}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\begin{cases}\mathcal{R}_{q}(\nu\mathbin{\|}\mu)-\frac{\log(1+t/C_{\mathsf{PI}}(\mu))}{q}&\text{if }\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\nu\mathbin{\|}\mu)}{(1+t/C_{\mathsf{PI}}(\mu))^{1/q}}&\text{if }\mathcal{R}_{q}(\nu\mathbin{\|}\mu)<1\,,\end{cases}$ (2.2)
@@
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M_{q}}{\varepsilon}\,,\quad\text{where }M_{q}:=\Bigl{\|}\frac{\mathrm{d}\pi_{0}^{X}}{\mathrm{d}\pi^{X}}\Bigr{\|}_{L^{q}(\pi^{X})}\,.$
We provide a version of this result in terms of $\chi^{q}$-divergence (which generalizes their $\chi^{2}$-result), as working with $\chi^{q}$-divergence leads to cleaner expressions and a bound on the contraction coefficient $\eta_{\chi^{q}}$ under ($\mathsf{PI}$). Our proof follows the argument sketched above, combining the general de Bruijn identity in terms of $\chi^{q}$-divergence under ($\mathsf{PI}$). Below, note that any log-concave probability measure satisfies ($\mathsf{PI}$) with $C_{\mathsf{PI}}<\infty$, as established in [KLS95, Bob99].
@@
Let $\mu$ be a log-concave probability measure over $\mathbb{R}^{d}$. For any $q\geq 2$, it holds that
@@
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\mu_{t}}\Bigl{[}\bigl{(}\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr{)}^{q}\,\bigl{|}\nabla\log\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr{|}^{2}\Bigr{]}\underset{\text{if }q\geq 2}{=}-\frac{q-1}{2}\,\Bigl{\|}\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\Bigr{\|}_{L^{q}(\mu_{t})}^{q}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,,$
@@
We combine general de Bruijn’s identity and the Rényi PI. For the former, we note that $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$ since ($\mathsf{PI}$) ensures exponential integrability [BU83, GM83] (i.e., $\int\exp(s\,|\cdot|)\,\mathrm{d}\mu<\infty$ if $s^{2}<4/C_{\mathsf{PI}}(\mu)$; see [BGL14, Proposition 4.4.2]). For the latter, note that $\nu_{t}$ and $\mu_{t}$ are smooth due to an elementary property of the heat kernel.
Using those two with $\chi^{q}+1=\lVert\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\rVert_{L^{q}(\mu_{t})}^{q}:=L^{q}\geq 1$, we deduce that
@@
Since $(q-1)\cdot x\,(1-x^{-1/(q-1)})\geq x-1$ for $x\geq 1$, we have
@@
Using $C_{\mathsf{PI}}(\mu_{t})\leq C_{\mathsf{PI}}(\mu)+t$ and solving the differential inequality, we obtain that
$\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\bigl{(}1+\frac{t}{C_{\mathsf{PI}}(\pi)}\bigr{)}^{-2/q}\,\chi^{q}(\nu\mathbin{\|}\mu)\,,$
@@
Under the same condition $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$, one can rederive (2.2) by following the argument in [CCSW22]. Using the identity $\mathcal{R}_{q}=\frac{1}{q-1}\,\log(1+\chi^{q})$ together with the chain rule and and de Bruijn’s identity, we obtain:
@@
When applying the Rényi PI, they used the inequality $21-e^{-x}\geq 1/$ 2 for $x\geq 1$ when $\mathcal{R}_{q}\geq 1$, and the inequality $21-e^{-x}\geq x/$ 2 for $x\leq 1$ when $\mathcal{R}_{q}\leq 1$. These inequalities are used to control the dissipation of the Rényi divergence in different regimes.
The mixing time of $\mathsf{PS}$ follows as a direct consequence of the contraction result above. Here, we ignore via the DPI additional contraction due to the backward step, although the backward step is also known to incur similar contraction under suitable regularity assumptions [CCSW22].
@@
We adapt this result to the setting of log-concave probability measures.
@@
For any $\varepsilon\in(0,1)$, $q\geq 2$, and log-concave probability measure $\pi^{X}$ over $\mathbb{R}^{d}$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
@@
Any log-concave probability measures satisfy ($\mathsf{PI}$) (and thus $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$). The claim then follows from the proposition above with $C_{\mathsf{PI}}(\pi^{X})\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\log d$ for log-concave probability measures [Kla23]. ∎
Correctness under algorithmic modifications.
It may not be immediately clear whether the transition kernel of $\mathsf{PS}$ remains unchanged under two algorithmic modifications: (1) setting a retry threshold $\tau$, and (2) restarting an iteration if all $\tau$ retries fail.
For completeness, we prove that the transition kernel is indeed preserved under these changes. We first show that imposing the threshold does not alter the kernel of the backward step.
@@
$=\frac{\sum_{i=1}^{\tau}\mathbb{P}(\text{fail in the first }(i-1)\text{ trials},\,X_{i}\ \text{accepted},\,X_{i}\in A)}{\sum_{i=1}^{\tau}\mathbb{P}(\text{success in the }i\text{-th trial})}$
@@
Restarting the forward step when $\tau$ retries fail also does not alter the transition kernel of $\mathsf{PS}$.
Lemma 2.7.
The transition kernel of $\mathsf{PS}$ is $P_{h}Q_{h}$.

### v2

@@ 2.1 Mixing analysis
The data-processing inequality (DPI) is a fundamental tool in information theory. For any probability measures $\mu,\nu$, Markov kernel $P$, $\varphi$-divergence $D_{\varphi}$, and $q\in(1,\infty]$, it holds that
@@
However, this alone is not sufficient to establish a mixing rate. It turns out that adding an independent Gaussian random variable to $X_{k}$ (i.e., convolving $\operatorname{law}X_{k}$ with $\gamma_{h}$) causes $\operatorname{law}X_{k}$ to contract directly toward $\pi^{X}*\gamma_{h}=\pi^{Y}$, with the contraction rate governed by $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$.
@@
where $\nu_{t}:=\nu*\gamma_{t}$ denotes the probability measure obtained by convolving $\nu$ with $\gamma_{t}$. Note that $0\leq\eta_{D}(\mu,t)\leq 1$ by the DPI.
Bounding $\eta_{D}$ relies on the well-known de Bruijn identity, $\partial_{t}\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{1}{2}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$, where the relative Fisher information is defined as $\mathsf{FI}(\nu\mathbin{\|}\mu):=\mathbb{E}_{\nu}[|\nabla\log\frac{\mathrm{d}\nu}{\mathrm{d}\mu}|^{2}]$; see rigorous proofs by Barron [BAR84] and Klartag and Ordentlich [KO25] (for its generalizations). Other derivations can be found in [GUO09, CCS+22], under the assumption that the probability measures involved are sufficiently smooth to justify differentiation under the integral sign and the vanishing of boundary terms in integration by parts.
Combining this identity with, for example, ($\mathsf{LSI}$) for $\mu$ (i.e., $\mathsf{KL}(\nu_{t}\mathbin{\|}\mu_{t})\leq\nicefrac{{C_{\mathsf{LSI}}(\mu_{t})}}{{2}}\,\mathsf{FI}(\nu_{t}\mathbin{\|}\mu_{t})$) and $C_{\mathsf{LSI}}(\mu_{t})\leq C_{\mathsf{LSI}}(\mu)+t$ [CHA04], one obtains a differential inequality in $\mathsf{KL}$; solving it yields
@@
Under similar regularity assumptions, the SDPI for Rényi divergence under ($\mathsf{PI}$), ($\mathsf{LSI}$), and Latała–Oleszkiewicz inequality was studied in [CCS+22] and further extended to constrained probability measures in [KVZ24]. These results were rigorously unified in [KO25] under a $\varphi$-Sobolev inequality [CHA04] (a generalization of ($\mathsf{PI}$) and ($\mathsf{LSI}$)), with regularity assumptions explicitly identified. We refer the reader to [KO25, Regularity conditions] for a detailed account.
Specifically, [CCS+22] showed that under the regularity assumptions, for any $q\geq 2$ and $\nu\ll\mu$, the following bounds hold:
[…]
@@
$N\lesssim qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M_{q}}{\varepsilon}\,,\quad\text{where }M_{q}=\Bigl\|\frac{\mathrm{d}\pi_{0}^{X}}{\mathrm{d}\pi^{X}}\Bigr\|_{L^{q}(\pi^{X})}\,.$
We provide a version of this result in terms of $\chi^{q}$-divergence (generalizing the $\chi^{2}$-result of [CCS+22]), as working with $\chi^{q}$-divergence yields cleaner expressions and a direct bound on the contraction coefficient $\eta_{\chi^{q}}$ under ($\mathsf{PI}$). Our proof follows the argument sketched above, combining the general de Bruijn identity in terms of $\chi^{q}$-divergence with ($\mathsf{PI}$). Below, we note that any logconcave probability measure satisfies ($\mathsf{PI}$) with $C_{\mathsf{PI}}<\infty$, as established in [KLS95, BOB99].
@@
Let $\mu$ be a logconcave probability measure over $\mathbb{R}^{d}$. For any $q\geq 2$, it holds that
@@
$\partial_{t}\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})=-\frac{q\,(q-1)}{2}\,\mathbb{E}_{\mu_{t}}\Bigl[\bigl(\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr)^{q}\,\bigl|\nabla\log\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\bigr|^{2}\Bigr]=-\frac{q-1}{2}\,\Bigl\|\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\Bigr\|_{L^{q}(\mu_{t})}^{q}\,\mathsf{RFI}_{q}(\nu_{t}\mathbin{\|}\mu_{t})\,$,
@@
We combine the general de Bruijn identity with the Rényi PI. For the former, we note that $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$ since ($\mathsf{PI}$) ensures exponential integrability [BU83, GM83] (i.e., $\int\exp(s\,|\cdot|)\,\mathrm{d}\mu<\infty$ for $s^{2}<4/C_{\mathsf{PI}}(\mu)$; see [BGL14, Proposition 4.4.2]). For the latter, $\nu_{t}$ and $\mu_{t}$ are smooth by an elementary property of the heat kernel.
Using $\chi^{q}+1=\lVert\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{t}}\rVert_{L^{q}(\mu_{t})}^{q}:=L^{q}\geq 1$ and combining the two, we deduce that
@@
Since $(q-1)\cdot x\,(1-x^{-1/(q-1)})\geq x-1$ for $x\geq 1$, we obtain
@@
Using $C_{\mathsf{PI}}(\mu_{t})\leq C_{\mathsf{PI}}(\mu)+t$ and solving the differential inequality yields
$\chi^{q}(\nu_{t}\mathbin{\|}\mu_{t})\leq\bigl(1+\frac{t}{C_{\mathsf{PI}}(\mu)}\bigr)^{-2/q}\,\chi^{q}(\nu\mathbin{\|}\mu)\,$,
@@
Under the same condition $\mathbb{E}_{\mu}[|\cdot|^{4}]<\infty$, one can rederive (2.2) by following the argument of [CCS+22]. Using the identity $\mathcal{R}_{q}=\frac{1}{q-1}\,\log(1+\chi^{q})$ together with the chain rule and de Bruijn’s identity, we obtain:
@@
When applying the Rényi PI, the inequality $1-e^{-x}\geq 1/2$ for $x\geq 1$ is used when $\mathcal{R}_{q}\geq 1$, and $1-e^{-x}\geq x/2$ for $x\leq 1$ when $\mathcal{R}_{q}\leq 1$. These control the dissipation of the Rényi divergence in each regime.
The mixing time of $\mathsf{PS}$ follows as a direct consequence of the contraction result above. We ignore via the DPI the additional contraction due to the backward step, although the backward step is also known to exhibit similar contraction under suitable regularity assumptions [CCS+22].
@@
We adapt this result to the setting of logconcave probability measures.
@@
For any $\varepsilon\in(0,1)$, $q\geq 2$, and logconcave probability measure $\pi^{X}$ over $\mathbb{R}^{d}$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
@@
Any logconcave probability measures satisfy ($\mathsf{PI}$) (and thus $\mathbb{E}_{\pi^{X}}[|\cdot|^{4}]<\infty$). The claim then follows from the proposition above with $C_{\mathsf{PI}}(\pi^{X})\lesssim\lVert\operatorname{cov}\pi^{X}\rVert\log d$ for logconcave probability measures [KLA23]. ∎
2.2 Per-step analysis
2.2.1 High-level idea of analysis
Let us first analyze $\mathsf{PS}_{\textup{unif}}$ as a one-attempt algorithm that may declare failure if the inner rejection sampler exceeds the threshold $\tau$. Fix an iteration horizon $N\in\mathbb{N}$, step size $h>0$, and inner cap $\tau\in\mathbb{N}$.
Let $S_{i}$ denote the survival event that failure has not occurred up to and including step $i$, with $S_{0}=\Omega$ and full success event $S:=S_{N}$. Let $C_{i}\geq 0$ be the number of membership queries used by the backward step at step $i$. The total query cost is then $C:=\sum_{i=1}^{N}\mathds{1}[S_{i-1}]\,C_{i}$. Note that
$\mathbb{E}C=\sum_{i=1}^{N}\mathbb{E}\bigl[\mathds{1}[S_{i-1}]\,C_{i}\bigr]\leq\sum_{i=1}^{N}\mathbb{E}C_{i}\,$,
so it suffices to bound $\mathbb{E}C_{i}$. Similarly, if $F_{i}$ denotes the event of failure at step $i$, and $q_{i}$ denotes the cap-hit probability at step $i$, then
$\mathbb{P}(F_{i})=\mathbb{E}\bigl[\mathds{1}[S_{i-1}]\,q_{i}\bigr]\leq\mathbb{E}q_{i},\qquad\mathbb{P}(S^{c})\leq\sum_{i=1}^{N}\mathbb{E}q_{i}(Y_{i})\,.$
These inequalities allow us to carry out the main per-step analysis under the clean unconditional intermediate laws (e.g., $Y_{i}\sim\pi_{i}^{Y}$), without explicitly tracking the conditional distributions at each step.
(1) Bias from failure.
It may not be immediately clear whether the transition kernel of $\mathsf{PS}$ is preserved under the algorithmic modification of imposing a retry threshold $\tau$. For completeness, we verify that the backward step is indeed unchanged by this modification.
@@ 2.2.1 High-level idea of analysis
[…]
@@
However, the law of output conditioned on success introduces a small bias, which is bounded roughly by the failure probability of the algorithm.
Lemma 2.7 (Bias from success, [KVZ26]).
Let $S$ be the event that the failure is not declared over the required $N$ iterations, and suppose $\mathbb{P}(S)\geq 1-\eta$. Let $\pi_{N}^{X}$ be the law of the uncapped $\mathsf{PS}_{\textup{unif}}$ output after $N$ iterations, and $\hat{\pi}_{N}^{X}$ be the law of $\mathsf{PS}_{\textup{unif}}$ conditioned on $S$ (i.e., $\hat{\pi}_{N}^{X}:=\operatorname{law}(X_{N}\mid S)$). Then,
$\frac{\mathrm{d}\hat{\pi}_{N}^{X}}{\mathrm{d}\pi_{N}^{X}}$ $\leq\frac{1}{1-\eta}\quad\textup{so}\quad\mathcal{R}_{\infty}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi_{N}^{X})\leq\log\frac{1}{1-\eta}\,$,
$\mathcal{R}_{q}(\hat{\pi}_{N}^{X}\mathbin{\|}\pi^{X})$ $\leq\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})+\frac{q}{q-1}\log\frac{1}{1-\eta}\qquad\text{for }q>1\,.$
(2) Restart-until-success.
A practical restart-until-success strategy repeats independent attempts until $S$ occurs. It outputs a sample with (conditional) law $\hat{\pi}_{N}^{X}$ and terminates almost surely whenever $\mathbb{P}(S)>0$. Moreover, if $\mathbb{P}(S)\geq 1-\eta$, then the expected number of attempts is at most $(1-\eta)^{-1}$ and
$\mathbb{E}[C\mid S]\leq\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,,\qquad\mathbb{E}C_{\text{restart}}=\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,.$
It therefore suffices to bound the one-attempt success probability $\mathbb{P}(S)$ and the unconditional expected cost $\mathbb{E}C$.
2.2.2 Per-step analysis
The analysis follows [KV25a] closely, which establishes the query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$. The only difference lies in the required number $N$ of iterations, namely the number needed for contraction in $\mathcal{R}_{q}$. Thus, we recall the relevant results from §2.1 of [KV25a] and adapt them minimally for our setting. With the parameter choices $Z=\frac{16NM_{2}}{\eta}$, $c=\frac{\log\log Z}{2\log Z}$, $\tau=Z^{2}\log^{3}Z$, and $h=\frac{c}{d^{2}}$, the failure probability satisfies $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$. The per-step complexity is bounded as follows: setting $q=1+\alpha$ with $\alpha\geq 1\vee\log\tau$, it follows that $\mathbb{E}_{\pi_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q}\log^{4}Z$.
We now prove the order-preserving complexity guarantees of $\mathsf{PS}_{\textup{unif}}$ under an $O(1)$-warm start and sufficient condition on $q$.
Proposition 2.8.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}\subset\mathcal{K}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to a membership oracle for $\mathcal{K}$. Given $\varepsilon>0$ , $\eta\in(0,1/2)$, $q\geq 2$, and an initial distribution $\pi_{0}$ with $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}$, initialize $\mathsf{PS}_{\textup{unif}}$ with $\pi_{0}$ and iterate it $N=\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(2d^{2}\log Z)^{-1}$ and $\tau=Z^{2}\log^{3}Z$ for $Z:=\frac{16NM_{2}}{\eta}$. If $q\geq 1+\log\tau$ and $M_{q}\leq 10$, then
  • The sufficient condition on $q$ is satisfied when $q\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{unif}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\pi)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\varepsilon}\log^{5}\frac{1}{\eta})$ membership queries in expectation.